# TRG-Net: An Interpretable and Controllable Rain Generator

### v1

@@ Abstract
Exploring and modeling rain generation mechanism is critical for augmenting paired data to ease training of rainy image processing models. Against this task, this study proposes a novel deep learning based rain generator, which fully takes the physical generation mechanism underlying rains into consideration and well encodes the learning of the fundamental rain factors (i.e., shape, orientation, length, width and sparsity) explicitly into the deep network. Its significance lies in that the generator not only elaborately design essential elements of the rain to simulate expected rains, like conventional artificial strategies, but also finely adapt to complicated and diverse practical rainy images, like deep learning methods. By rationally adopting filter parameterization technique, we first time achieve a deep network that is finely controllable with respect to rain factors and able to learn the distribution of these factors purely from data. Our unpaired generation experiments demonstrate that the rain generated by the proposed rain generator is not only of higher quality, but also more effective for deraining and downstream tasks compared to current state-of-the-art rain generation methods. Besides, the paired data augmentation experiments, including both in-distribution and out-of-distribution (OOD), further validate the diversity of samples generated by our model for in-distribution deraining and OOD generalization tasks.
keywords:
Rain generation, Interpretable network, Unpaired data generation, Data augmentation
1 Introduction
@@
Exploring and modeling rain generation mechanism is critical for augmenting paired data to ease the training of rainy image processing models. This is especially meaningful for current deep learning based image rain removal methods, where the effectiveness is highly dependent on paired rainy and rainless images [1, 2, 3, 4]. However, ideal data pairs are always hard to collect, particularly in the complicated and diverse rain scenarios in practice. Besides, the rain generation task is also of great potential value in revealing insightful characteristics underlying real rains, helpful in guiding sound policies for dealing with practical severe weather emergencies [5, 6, 7].
@@
The early methods for investigating how to generate rainy images are mainly physical rendering based artificial synthesis manners [8, 9, 10]. Specifically, as shown in Fig. 1, there are mainly three parts for generating a rainy image. The first part is the rain kernel model, for generating convolution kernels that involve the orientation, length, width and other shape information of the rain streaks. The second part is the rain map model, which is decided by the sparsity and position of the rains. By convolving the rain kernel and rain map, we can obtain a lifelike rain streak image. The final part is a merging model that simulates the interaction between rain streaks and a background image (e.g., rain streak photometry and fog-like rain). By carefully designing the three generation parts, rainy images can be synthesized with controllable shape, orientation, width, length and sparsity, even without any model training process. This manner has been widely used in the generation of rain scenes in movies, video games, animations and artistic works [11, 12].
Although this category of methods can synthesize visually lifelike rainy images in the sense of human intuition, it still lacks realism to use the synthesized rainy/rainless images to train deep neural networks for deraining [13, 14]. This is because the generation models of these methods are designed purely by human subjectivity, which inevitably leads to differences with the more complicated and diverse rains in real scenes.
@@
Very recently, deep learning (DL) based manners have been exploited for this task. The basic idea is to adopt generative adversarial networks (GANs) to train a rain generator and a discriminator, where the generator captures the rainy data distribution, and the discriminator distinguishes the difference between the real rains and the generated ones [15, 16, 17]. In this way, the generated rains are forced to get close to the real rains in terms of the probability distribution through the training with real samples. It has been shown that the rainy/rainless image pairs generated in this way can help to achieve a great performance improvement in deraining tasks [18, 19, 20].
@@
However, the current rain synthesis methods based on GANs still have obvious drawbacks. The most critical one is the design of rain generators. In particular, the current rain generators are usually assembled with some off-the-shelf network modules in commonly-adopted deep learning toolkits, e.g., convolutional neural networks (CNNs), which is indeed a “black box” without considering the intrinsic rain generation mechanism, and thus increases the learning difficulty and data requirements. Moreover, this insufficient knowledge modeling issue makes these methods hardly be effectively trained under the condition of inadequate paired training data, like those with unpaired rainy/rainless images [18, 19, 20, 21, 22]. A more popular way is to use these GAN-based generators just for data augmentation from a small number of paired data samples [18, 19]. However, the problem of paired data collection still remains in this manner. It is thus critical to develop a rain generator capable of more faithfully delivering real rain generation mechanism for rain generation, with higher generation capability and lower data requirements.
@@
Fortunately, the generation mechanism of rain streaks has been well studied in the previous physical rendering based artificial synthesis methods, where one can find that rain streaks are mainly decided by their shapes, orientations, widths, lengths and sparsity [8, 9, 23], as shown in Fig. 1. In this paper, we focus on intrinsically combining the advantages of the physical rendering based artificial synthesis methodology and the deep learning based one to design a dedicated generator for rain generation.
@@
Specifically, we aim to design a deep learning based rain generator which is controllable with respect to the rain factors (i.e., shape, orientation, width, length of the rain kernel and sparsity of the rain map) just like conventional artificial synthesis methods, while the distribution of these rain factors can be learned through the generative adversarial manner instead of being manually preset. This is actually not that trivial to achieve, since the current CNNs cannot be easily controlled by the aforementioned rain factors. For example, for controlling the orientation of rain streak generation, we have to dynamically rotate an entire CNN architecture in its spatial dimensions, which is not easy to implement with common CNN architecture. It is even more difficult to learn the orientation distribution of the rain streak in the whole rainy image dataset. In this paper, we exploit the filter parametrization technique and firstly design such an expected rain factor transformable convolution network for rain generation, where the aforementioned goals can all be finely achieved. Specifically, the main contributions of this paper are summarized as follows:
@@
  • We construct a new rain generator, which can properly extract fundamental rain factors underlying rainy images, including shape, orientation, width, length and sparsity, directly from data in a purely automatic manner. Its significance lies in that the model can both elaborately design fundamental elements to simulate expected rains like conventional artificial rendering manners, and finely adapt to complicated and diverse practical rain forms like recent deep learning methods. To the best of our knowledge, this should be the first method that is not only controllable with respect to rain factors, but also can learn the distribution of these factors purely from data. (e.g., Fig. 2 shows the extracted distribution about the orientation factor from Rain100L, finely complying with its real situation). The model is thus with intrinsic interpretability and essential controllability.
  • To alleviate the difficulty in embedding controllable and learnable rain factors into deep networks, we construct a transformable convolution framework. Specifically, we exploit the filter parametrization technique [24] for representing convolution filters to ameliorate the original discrete expression as continuous when constructing CNNs. This makes the expected transformation operators, such as rotation and scaling, able to be imposed on the convolution filters more easily and soundly, while keeping them learnable. Under such transformable convolutions, we can naturally construct a transformable rain kernel model (in Sec. 4.2.1) and rotatable CNNs (in Secs. 4.2.2 and 4.2.3). Moreover, we present a new total variation (TV) regularizer, called rotatable TV regularizer (in Sec. 4.3), capable of adaptively adjusting the orientation for calculating variations and adopting higher penalty along the rain streak orientation. The aforementioned functions could be potentially applied to a wider range of tasks.
@@
The paper is organized as follows. Sec. 2 reviews necessary related works. Sec. 3 presents the transformable convolution framework employed in our method. Sec. 4 then proposes details of our proposed transformable rain generator. Sec. 5 demonstrates comprehensive experiments to substantiate the effectiveness of the proposed method. The paper is finally concluded with future work.
2 Related Work
2.1 Rain Synthesis Methods
@@
The most straightforward way to collect rainy/rainless image pairs is to utilize rainy videos, where rainy images come from the video frames and rainless images are estimated from a sequence of video frames [25, 26]. While this method is able to collect the rainy images in real scenes, the resulting rainless images often contain remnants of rain streaks or lost image details.
@@
Physical rendering based artificial synthesis is a common method that synthesizes rainy/rainless image pairs. The appearance of rain was studied early based on a raindrop oscillation model by Garg and Nayar [8, 9]. The rain streaks are considered to be the motion blur of the raindrops and can be synthesized with controllable shape, orientation, width, length and sparsity, by carefully setting the parameters of the rain model. Many rain datasets are constructed using the photorealistic rendering technique [^1] https://www.photoshopessentials.com/photo-effects/rain/. In general, rain streaks are first synthesized with physical models [8, 9], and then the rain streaks are blended to rainless background images to get the paired rainy images. Some works assume that rain streaks are simply superimposed to the background images, such as Rain100L [7, 27] and Rain800 [28]. Some works, such as Rain100H [7], Cityscapes [29, 30] and Kitti [29, 30], try to explore a more complex merge relationship between rain layer and background image. These large-scale paired rain datasets offer the training set for deraining networks and can quantitatively measure the effectiveness of deraining models. However, the rain streaks synthesized by these methods often tend to deviate from the real rains with much more complicated and diverse configurations. Such deviation usually leads to a degradation of the performance of the deraining models when dealing with the real rainy images.
@@
Recently, GANs [15, 16] are employed for the rain generation where the probability distribution of the generated rain is forced to get close to that of the real rain through the training of certain generation network, which has achieved great success. Typically, [29, 30, 31] generated rains using the manually-designed rain models, and GANs are only adopted for learning the wetness scene. [18, 19, 20] generated rains using a generator stacked from some existing network modules and ignored the intrinsic rain generation mechanism, which is indeed a “black box” network. These methods tend to increase the learning difficulty and data requirements due to the neglecting of rain generation mechanism. Moreover, this insufficient knowledge modeling issue also makes these methods hardly able to be sufficiently trained on the condition of inadequate paired training data, like those with unpaired rainy/rainless images [18, 19].
Different from previous rain synthesis methods, our aim is to intrinsically combine the advantages of both physical based methods and deep learning based methods, and design a deep rain factor transformable rain generator with intrinsic interpretability and essential controllability.
2.2 Single Image Rain Removal
@@
There are some works trying to tackle this issue in a semi-supervised [13, 14] or unsupervised way [35, 22], but their performance is usually limited due to the lack of effective supervision information. Therefore, it is critical to design a rain generator for faithfully simulating real rain data with essentially more complexity and flexibly constructing paired datasets.
3 Transformable Convolution Framework
Traditional convolution kernels are usually discrete, which can’t be arbitrarily and accurately rotated or scaled. It is then natural to exploit continuous representation of convolution kernels for easily constructing transformable convolution. To this end, we employ filter parametrization methods to construct the overall framework of the transformable convolution as is shown in Fig. 3.
Specifically, the filter parametrization approach [36] regards a discrete convolution kernel ${\phi}\in\mathbb{R}^{p\times p}$ as the discretization of a 2D continuous function $\tilde{\phi}$ on the $\left[-\nicefrac{{(p-1)}}{{2}},\nicefrac{{(p-1)}}{{2}}\right]^{2}$ area. A typical formulation for constructing $\tilde{\phi}(x)$ is [37]:
@@ 3 Transformable Convolution Framework
The key issue here is the choice of the basis function set. Especially, for rain generation task which requires pixel-level accuracy, it is important to choose a basis function set so that any discrete filter can be accurately represented, i.e., for any ${\phi}$, there exists a set of $w_{n}$ so that $\tilde{\phi}(x_{ij})={\phi}_{ij}$. Besides, another issue here is to avoid aliasing effect [^2] The aliasing effect here is caused by the insufficient sampling rate of discrete filter when the bases frequency is too high, resulting in an incorrect transformation result. For a detailed analysis of the aliasing effect please refer to [24]., otherwise the convolution kernel can be very unstable when being transformed [36, 37, 24].
@@
It should be noted that $w_{n},n=1,\cdots,N$ are learnable parameters when utilizing this convolution kernel in deep CNNs.
@@
where $\theta,s_{l},s_{w}$ are the rotation degree, the length parameter, and the width parameter, respectively. It should be noted that $s_{l}$ and $s_{w}$ are actually inversely proportional to the length and width, respectively, since $T$ is an inverse transformation matrix. In addition, we can find that all parameters (including transform parameters $\Omega$ and coefficient parameters $w_{n}$) of the convolution operation defined in Eq. (6) are easy to implement gradient backpropagation in deep learning frameworks. Therefore, the proposed transformable convolution is potentially applicable to a wider range of the architecture designs in deep networks.
4 Transformable Rain Generator

### v2

@@ Abstract
Exploring and modeling rain generation mechanism is critical for augmenting paired data to ease training of rainy image processing models. Most of the conventional methods handle this task in an artificial physical rendering manner, through elaborately designing fundamental elements constituting rains. This kind of methods, however, are over-dependent on human subjectivity, which limits their adaptability to real rains. In contrast, recent deep learning methods have achieved great success by training a neural network based generator from pre-collected rainy image data. However, current methods usually design the generator in a “black-box” manner, increasing the learning difficulty and data requirements. To address these issues, this study proposes a novel deep learning based rain generator, which fully takes the physical generation mechanism underlying rains into consideration and well encodes the learning of the fundamental rain factors (i.e., shape, orientation, length, width and sparsity) explicitly into the deep network. Its significance lies in that the generator not only elaborately design essential elements of the rain to simulate expected rains, like conventional artificial strategies, but also finely adapt to complicated and diverse practical rainy images, like deep learning methods. By rationally adopting filter parameterization technique, the proposed rain generator is finely controllable with respect to rain factors and able to learn the distribution of these factors purely from data without the need for rain factor labels. Our unpaired generation experiments demonstrate that the rain generated by the proposed rain generator is not only of higher quality, but also more effective for deraining and downstream tasks compared to current state-of-the-art rain generation methods. Besides, the paired data augmentation experiments, including both in-distribution and out-of-distribution (OOD), further validate the diversity of samples generated by our model for in-distribution deraining and OOD generalization tasks. Code is available at https://github.com/pzq-xjtu/TRG-Net.
Index Terms:
Rain generation, Interpretable network, Unpaired data generation, Data augmentation.
I Introduction
@@
Exploring and modeling rain generation mechanism is critical for augmenting paired data to ease the training of rainy image processing models. This is especially meaningful for current deep learning based image rain removal methods, where the effectiveness is highly dependent on paired rainy and rainless images [1, 2, 3, 4]. However, ideal data pairs are always hard to collect, particularly in the complicated and diverse rain scenarios in practice. Besides, the rain generation task is also of great potential value in revealing insightful characteristics underlying real rains, helpful in guiding sound policies for dealing with severe weather emergencies [5, 6, 7].
@@
The early methods for investigating how to generate rainy images are mainly physical rendering based artificial synthesis manners [8, 9, 10]. The physical rendering methods model rain streaks, the main rain appearance in rainy images, as the motion blur resulting from raindrops under the influence of gravity and wind, where the raindrops are characterised by an elaborate oscillation model. Specifically, as shown in Fig. 1, there are mainly three parts for generating a rainy image. The first part is the rain kernel model, for generating convolution kernels that involve the orientation, length, width and other shape information of the rain streaks. The second part is the rain map model, which is decided by the sparsity and position of the rains. By convolving the rain kernel and rain map, we can obtain a lifelike rain streak image. The final part is a merging model that simulates the interaction between rain streaks and a background image (e.g., the fog-like rain formed by rain accumulation in 3D space). By carefully designing the three generation parts, rainy images can be synthesized with controllable shape, orientation, width, length and sparsity, even without any model training process. This manner has been widely used in the generation of rain scenes in movies, video games, animations and artistic works [11, 12].
@@
Although this category of methods can synthesize visually lifelike rainy images in the sense of human intuition, it still lacks realism to use the synthesized rainy/rainless images to train deep neural networks for deraining [13, 14]. This is because the rain factors (e.g., orientation, width, length and sparsity) of these methods are designed purely by human subjectivity, which inevitably leads to differences with the more complicated and diverse rains in real scenes.
@@
Recently, deep learning (DL) based manners have been exploited for this task. The basic idea is to adopt generative adversarial networks (GANs) to train a rain generator and a discriminator, where the generator captures the rainy data distribution, and the discriminator distinguishes the difference between the real rains and the generated ones [15, 16, 17]. In this way, the generated rains are forced to get close to the real rains in terms of the probability distribution through the training with real samples. It has been shown that the rainy/rainless image pairs generated in this way can help to achieve a great performance improvement in deraining tasks [18, 19, 20, 21].
@@
However, the current rain synthesis methods based on GANs still have obvious drawbacks. The most critical one is the design of rain generators. In particular, the current rain generators are usually assembled with some off-the-shelf network modules in commonly-adopted deep learning toolkits, e.g., convolutional neural networks (CNNs), which is indeed a “black box” without considering the intrinsic rain generation mechanism, and thus increases the learning difficulty and data requirements. Moreover, this insufficient knowledge modeling issue makes these methods hardly be effectively trained under the condition of inadequate paired training data, like those with unpaired rainy/rainless images [18, 19, 20, 22, 23]. It is thus critical to develop a rain generator capable of more faithfully delivering real rain generation mechanism for rain generation, with higher generation capability and lower data requirements.
@@
Fortunately, the generation mechanism of rain streaks has been well studied in the previous physical rendering based artificial synthesis methods, where one can find that rain streaks are mainly decided by their shapes, orientations, widths, lengths and sparsity [8, 9, 24], as shown in Fig. 1. Significant progress has been made in a very recent study [21], which incorporates these rain factors into deep networks and learns the connection between rain factors and deep networks using additional labels for rain factors. However, the requirement for these rain factor labels would further enhance the difficulty in data collecting.
@@
In this paper, we focus on more intrinsically embedding the physical model of rain generation into deep networks to design a dedicated generator for rain generation, which can be controlled with the rain factors and extract the distribution of rain factors without additional data labels on rain factors. Specifically, we aim to mathematically model the relationship between rain factors (i.e., shape, orientation, width, length of the rain kernel and sparsity of the rain map) and CNN network modules, while maintaining the learnable characteristics of CNN network. Since the calculations to rain factors are designed manually, there is thus no need for additional rain factor labels in the training process. This is actually not trivial to achieve. For example, for controlling the orientation of rain streak generation, we have to dynamically rotate an entire CNN architecture in its spatial dimensions, which is not easy to implement with common CNN architecture. It is even more difficult to learn the orientation distribution of the rain streak in the whole rainy image dataset.
To alleviate the above-mentioned difficulty, we exploit the filter parametrization technique and firstly design such an expected rain factor transformable convolution network for rain generation, where the aforementioned goals can all be finely achieved. The main contributions of this paper are summarized as follows:
@@
  1. 1.
    We construct a new rain generator, which can properly extract fundamental rain factors underlying rainy images, including shape, orientation, width, length and sparsity, directly from data in a purely automatic manner. Its significance lies in that the model can both elaborately design fundamental elements to simulate expected rains like conventional artificial rendering manners, and finely adapt to complicated and diverse practical rain forms like recent deep learning methods. (e.g., Fig. 2 shows the extracted distribution about the orientation factor from Rain100L, finely complying with its real situation). The model is thus with intrinsic interpretability and essential controllability, which can be trained without requiring data labels about the rain factors.
  2. 2.
    To alleviate the difficulty in embedding controllable and learnable rain factors into deep networks, we construct a transformable convolution framework. Specifically, we exploit the filter parametrization technique [25] for representing convolution filters to ameliorate the original discrete expression as continuous when constructing CNNs. This makes the expected transformation operators, such as rotation and scaling, able to be imposed on the convolution filters more easily and soundly, while keeping them learnable. Under such transformable convolutions, we can naturally construct a transformable rain kernel model (in Sec. IV-B1) and rotatable CNNs (in Secs. IV-B2 and IV-B3). Moreover, we present a new total variation (TV) regularizer, called rotatable TV regularizer (in Sec. IV-C), capable of adaptively adjusting the orientation for calculating variations and adopting higher penalty along the rain streak orientation. The aforementioned functions could be potentially applied to a wider range of tasks.
  3. 3.
@@
The paper is organized as follows. Sec. II reviews necessary related works. Sec. III presents the transformable convolution framework for alleviating the difficulty that encodes controllable and learnable rain factors into deep networks. Sec. IV then proposes details of our proposed transformable rain generator. Sec. V demonstrates comprehensive experiments to substantiate the effectiveness of the proposed method. The paper is finally concluded with future work.
II Related Work
II-A Rain Synthesis Methods
@@
The most straightforward way to collect rainy/rainless image pairs is to utilize rainy videos, where rainy images come from the video frames and rainless images are estimated from a sequence of video frames [26, 27, 28]. While this method is able to collect large-scale rainy images in real scenes, the resulting rainless images often contain remnants of rain streaks or lost image details.
@@
Physical rendering based artificial synthesis is a common method that synthesizes rainy/rainless image pairs. The appearance of rain was studied early based on a raindrop oscillation model by Garg and Nayar [8, 9]. The rain streaks, the main appearance of rain in rainy images, are considered to be the motion blur of the raindrops and can be synthesized with controllable shape, orientation, width, length and sparsity, by carefully setting the parameters of the rain model. Many large-scale rain datasets are constructed using the photorealistic rendering technique [^1] https://www.photoshopessentials.com/photo-effects/rain/, such as Rain100L [7], Rain100H [7], Cityscapes [29] and Kitti [29]. These large-scale paired rain datasets offer the training set for deraining networks and can quantitatively measure the effectiveness of deraining models. However, the rain streaks synthesized by these methods often tend to deviate from the real rains with much more complicated and diverse configurations. Such deviation usually leads to a degradation of the performance of the deraining models when dealing with the real rainy images.
@@
Recently, GANs [15, 16] are employed for the rain generation where the probability distribution of the generated rain is forced to get close to that of the real rain through the training of certain generation network, which has achieved great success. Typically, [18, 19, 20] generated rains using a generator stacked from some existing network modules and ignored the intrinsic rain generation mechanism, which is indeed a “black box” network. These methods tend to increase the learning difficulty and data requirements due to the neglecting of rain generation mechanism. Moreover, this insufficient knowledge modeling issue also makes these methods hardly able to be sufficiently trained on the condition of inadequate paired training data, like those with unpaired rainy/rainless images [18, 19]. Very recent work by Yu et al. [21] has been made significant progress in this task, which incorporates rain factors into deep networks using extra labels of these rain factors. However, the requirement for these rain factor labels would further enhance the difficulty in data collecting.
Different from previous rain synthesis methods, we focus on more intrinsically embedding the physical model of rain generation into deep networks to design a dedicated generator for rain generation, which can be controlled with the rain factors and extract the distribution of rain factors without the need for additional data labels on rain factors.
II-B Single Image Rain Removal
@@
There are some works trying to tackle this issue in a semi-supervised [13, 14] or unsupervised way [33, 23], but their performance is usually limited due to the lack of effective supervision information. Therefore, it is critical to design a rain generator for faithfully simulating real rain data with essentially more flexibly constructing paired datasets.
III Transformable Convolution Framework
In this section, we propose a transformable convolution framework to alleviate the difficulty in embedding controllable and learnable rain factors into deep networks. Traditional convolution kernels in deep networks are usually discrete, which can’t be arbitrarily and accurately rotated or scaled. It is then natural to exploit continuous representation of convolution kernels for easily constructing transformable convolution. To this end, we employ filter parametrization methods to construct the overall framework of the transformable convolution, as shown in Fig. 3.
Specifically, the filter parametrization approach [34] regards a discrete convolution kernel ${\phi}\in\mathbb{R}^{p\times p}$ as the discretization of a 2D continuous function $\tilde{\phi}$ on the $\left[-\nicefrac{{(p-1)}}{{2}},\nicefrac{{(p-1)}}{{2}}\right]^{2}$ area. A typical formulation for constructing $\tilde{\phi}(x)$ is:
@@ III Transformable Convolution Framework
The key issue here is the choice of the basis function set. Especially, for rain generation task which requires pixel-level accuracy, it is important to choose a basis function set so that any discrete filter can be accurately represented, i.e., for any ${\phi}$, there exists a set of $w_{n}$ so that $\tilde{\phi}(x_{ij})={\phi}_{ij}$. Besides, another issue here is to avoid aliasing effect [^2] The aliasing effect here is caused by the insufficient sampling rate of discrete filter when the bases frequency is too high, resulting in an incorrect transformation result. For a detailed analysis of the aliasing effect please refer to [25]., otherwise the kernel can be very unstable when being transformed [25].
@@
It should be noted that $w_{n},n=1,\cdots,N$ are learnable parameters when utilizing this convolution kernel in the convolution layers of deep networks.
@@
where $\theta,s_{l},s_{w}$ are the rotation degree, the length parameter, and the width parameter, respectively. It should be noted that $s_{l}$ and $s_{w}$ are actually inversely proportional to the length and width, respectively, since $T_{\Omega}$ is an inverse transformation matrix. In addition, we can find that all parameters (including transform parameters $\Omega$ and coefficient parameters $w_{n}$) of the convolution operation defined in Eq. (6) are easy to implement gradient backpropagation in deep learning frameworks. Therefore, the proposed transformable convolution should be potentially applicable to a wider range of the architecture designs in deep networks. In the following section, we employ the proposed transformable convolution framework to design a transformable rain generator that encodes the controllable and learnable rain factors.
IV Transformable Rain Generator